OpenStudy (anonymous):

How do you evaluate arccsc 2?

OpenStudy (anonymous):

arccsc 2 has the same meaning as 1/ sin x = 2 sinx =1/2 just keep in mind arc csc and arc sin they have same range so we will not lose any answer if sinx =1/2 then arc sin (1/2) =? arc sin (1/2) = 30 degree or Pi/6